OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? –2x – 11y = 5 y= 2/11x +5/11; slope: 5/11; y-intercept: 2/11 y= 2/11x +5/11; slope: 2/11; y-intercept: 5/11 y= 2/11x - 5/11; slope: 2/11; y-intercept: -5/11 y= -2/11x - 5/11; slope: -2/11; y-intercept: -5/11

OpenStudy (anonymous):

The last one I believe

OpenStudy (anonymous):

I really need to get this rite cause I need a good grade. lol

OpenStudy (anonymous):

I'd be willing to bet money on it.. It's the last one. Do you need to know how? I can explain. But if you only need to know the answer it's D

OpenStudy (anonymous):

No , its fine. Thanks .

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

-2x-11y=5 -11y = 2x + 5 y = (-11/2)x - 5/11 slope -11/2 yintercept -5/11

OpenStudy (anonymous):

Precalc_sucks is right

OpenStudy (anonymous):

thanks yall .